the length of a rectangle exceeds its width by three inches the area of the rectangle is seventy square inches
1
Expert's answer
2012-10-10T11:19:20-0400
Let's denote the length of the rectangle by L and it's width by W. Then we have the following system of equations:

L = W + 3, (1)
L*W = 70. (2)

Let's solve it.

(2): (W + 3)*W = 70 <==> W² + 3W - 70 = 0 ==> W = 7;
&
L = 7 + 3 = 10.

So, the length of the rectangle is 10 inches and the width is 7 inches.